At what temperature will both celius & Fahrenheit scale read the same value?

Grade:11

1 Answers

Arun
5 years ago

We know the relation between °F and °C scale

°F = (9/5)°C + 32................(i)

Let the temperature will be x at which both the celsius and fahrenheit scales read the same value.

Putting x in place of °F and °C scale in (i) equation:

x = 9/5 (x) + 32

x - 32 = 9x/5

5x - 160 = 9x

4x = -160

x = -40

So at -40°, both celsius and fahrenheit scales read the same value.